# Detecting Multimedia Generated by Large AI Models: A Survey

### v1

@@ 6 Conclusion
This paper provides the first systematic and comprehensive survey covering existing research on detecting multimedia from text, image, audio, and video to multimodal content generated by large AI models. We introduce a novel taxonomy for detection methods within each modality, categorizing them under two primary frameworks: pure detection (focusing on improving detection accuracy) and beyond detection (integrating attributes like generalizability, robustness, and interoperability to detectors). Additionally, we have outlined the sources contributing to detection, such as generation mechanisms of LAIMs, public datasets, and online tools. Finally, we pinpoint current challenges in this field and propose potential directions for future research. We believe that this survey serves as the initial contribution to addressing a notable academic gap in this field, aligning with global AI security initiatives, thereby upholding the authenticity and integrity of digital information.

### v4

@@ 6 Conclusion
This paper provides the first systematic and comprehensive survey covering existing research on detecting multimedia from text, image, audio, and video to multimodal content generated by large AI models. We introduce a novel taxonomy for detection methods within each modality, categorizing them under two primary frameworks: pure detection (focusing on improving detection accuracy) and beyond detection (integrating attributes like generalizability, robustness, and interoperability to detectors). We further analyze them with modality-specific insights and recommendations. Additionally, we have outlined the sources contributing to detection, such as generation mechanisms of LAIMs, public datasets, online tools, evaluation metrics. We also emphasize the importance of detecting social media processed multimedia throughout the survey, systematically reviewing social media specific datasets, detection methods, and unique challenges across modalities. Finally, we pinpoint current challenges in this field and propose potential directions for future research in a modality-specific manner. We believe that this survey serves as the initial contribution to addressing a notable academic gap in this field, aligning with global AI security initiatives, thereby upholding the authenticity and integrity of digital information.
Acknowledgments
This work is supported by the U.S. National Science Foundation (NSF) under grant IIS-2434967 and the National Artificial Intelligence Research Resource (NAIRR) Pilot and TACC Lonestar6. The views, opinions and/or findings expressed are those of the author and should not be interpreted as representing the official views or policies of NSF and NAIRR Pilot.